OpenStudy (anonymous):

What is the value of b2 - 4ac for the following equation? x(x + 8) = 9 a) 28 b) 64 c) 100

OpenStudy (anonymous):

first simplify x(x + 8) = 9 x^2 + 8x = 9 x^2 + 8x - 9 = 0 a = 1 b = 8 and c = -9 b^2 - 4ac 8^2 - 4(1)(-9) 64 + 36 100